OpenStudy (anonymous):

4/21+3/21 show all work and answer in simpliest form

OpenStudy (anonymous):

$$ \frac{4}{21} + \frac{3}{21} = \frac{7}{21} = \frac{7\times1}{7\times3} = \frac{1}{3} $$